The R-Bar-M Ranch in Montana would like a new mechanized barn, which will require a $600,000...

  1. The R-Bar-M Ranch in Montana would like a new mechanized barn, which will require a $600,000 initial cash outlay. The barn is expected to provide after-tax annual cash savings of $90,000 indefinitely (for practical purposes of computation, forever). The ranch, which is incorporated and has a public market for its stock, has a weighted average cost of capital of 14.5 percent. For this project, Mark O. Witz, the president, intends to provide $200,000 from a new debt issue and another $200,000 from a new issue of common stock. The balance of the financing would be provided internally by retaining earnings. The present value of the after-tax flotation costs on the debt issue amount to 2 percent of the total debt raised, whereas flotation costs on the new common stock issue come to 15 percent of the issue. What is the net present value of the project after allowance for flotation costs? Should the ranch invest in the new barn?
Answer #1

Answer:

Out flows:

Initial outlay = $600,000

Flotation cost:

Debt after tax flotation cost in % = 2%

Debt flotation cost = 200000 / (100% - 2%) * 2% = $4081.633

Equity flotation cost in % = 15%

Equity flotation cost = 200000 / (100% - 15%) * 15% =$35,294.118

Total flotation costs = 4081.633 + 35294.118 = $39375.751

Inflows:

Expected to generate perpetual after tax annual cash savings = $90,000

WACC = 14.5%

Present Value of inflows = Perpetual after tax annual cash savings / WACC = 90000 / 14.5% = $620,689.655

NPV:

Net present value of the project after allowance for flotation costs = 620689.655 - 600,000 - 39375.751 = - $18,686.10

Net present value of the project after allowance for flotation costs = - $18,686.10

The ranch should not invest in the new barn. The NPV is negative.